Integrate with respect to x: ∫ xdx

a. x^2/2+c

b. 5/8

c. (-8)/5

d. 8/5
1
Expert's answer
2013-05-24T08:36:16-0400
int; x dx this is one of table integrals. As we know,& ∫ x^n dx=x^(n+1)/(n+1)+c
So in our case we obtain& ∫ x dx=x^(1+1)/(1+1)+c=x^2/2+c
Thus, the right answer is a.